There are balls numbered 1 through 7 in a box. Two balls are selected at random in succession without replacement, and the number on each ball is noted.
What probability should be assigned to the event that at least one ball has an odd number?
How do I set this up?

Answer by stanbon(75887)

There are balls numbered 1 through 7 in a box. Two balls are selected at random in succession without replacement, and the number on each ball is noted.
What probability should be assigned to the event that at least one ball has an odd number?
---------
Number of odd-numbered balls: 4
---
P(2 odd when 2 selected without replacement) = 4C2/7C2 = (4*3)/(7*6)
= 12/42
= 6/21
= 2/7
============
Cheers,
Stan H.